OpenStudy (anonymous):

find the greatest commom factor of 2730 and 9350.

OpenStudy (cwrw238):

thats a long one - its best to you use euclids method here - can't remember it!

OpenStudy (anonymous):

a=bq+r.is it?

OpenStudy (anonymous):

Prime Factorizations: 2730 = 2*3*5*7*13 9350 = 2 * 5 *5 * 11 * 17 GCD is therefore 10. :-) Is that one of your choices?

OpenStudy (anonymous):

ya......thx

OpenStudy (cwrw238):

don't know i think you subtract them first 9350 - 2730 = 6620 then 6620 - 2730 = 3890 3890 - 2730 = 1160 2730 - 1160 = 1570 1570 - 1160 = 410 1160 - 410 = 750 750-410 = 340 410 - 340 = 70 gcf of 70 and 340 70 = 2 * 5 * 7 340 = 2 * 2 *5 * 13 gcf = 5 * 2 = 10

OpenStudy (cwrw238):

lol - it would have been easier to factor!!!!

OpenStudy (anonymous):

That's what I did ;-) Prime factorization. /me points

OpenStudy (anonymous):

ya by lemma its easy ..anyways thx

OpenStudy (cwrw238):

yep - you really need to do the euclid thing on a spreadsheet

OpenStudy (anonymous):

kk.

OpenStudy (cwrw238):

when you have really high numbers then euclid is very useful

OpenStudy (anonymous):

ya it will:)